# L’approximation arithmétique de Fujita sur courbes adéliques

### v2

@@ 3.4 Graded linear series over trivially valued fields
Let $\overline{E}$ be an adelic vector bundle over an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, and $\left(\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E})\right)_{t\in\mathbb{R}}$ be the Harder-Narasimhan filtration of $\overline{E}$ defined in Definition 2.2.3. We define the norm .$\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}$ on $E$ over $(K,|\raisebox{1.72218pt}{.}|_{0})$ as
.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}:&E&\longrightarrow&\mathbb{R}^{+}\\ &s&\mapsto&\exp\left(-\sup\{t\in\mathbb{R}\mid s\in\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E})\}\right).\end{array}$ (7)
@@
Let $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ be a normed graded linear series over an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$. For every $\overline{E}_{D}$, we denote by .$\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}}$ the norm (7) on $E_{D}$ indexed by $D\in\mathbb{N}$. Then $(E_{D},\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}})$ an adelic vector bundle over $S_{0}$.
@@
$\|s\cdot t\|^{\operatorname{HN}}_{m+n}\leqslant\exp(f(m)+f(n))\|s\|^{\operatorname{HN}}_{m}\|t\|^{\operatorname{HN}}_{n},$
@@
In order to obtain a new family of norms which are sub-multiplicative, we introduce the spectral norm .$\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ of .$\|\raisebox{1.72218pt}{.}\|^{\operatorname{HN}}_{D}$ as
.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}:&H^{0}(X,L^{\otimes D})&\longrightarrow&\mathbb{R}_{\geqslant 0}\\ &s&\mapsto&\lim\limits_{n\rightarrow\infty}\left(\|s^{n}\|_{nD}^{\operatorname{HN}}\right)^{1/n}.\end{array}$ (8)
@@
We denote by $\mathcal{F}_{\mathrm{sp}}$ the $\mathbb{R}$-filtration on $\overline{E}_{\operatorname{HN},D}$ induced by the filtrated structure $\lambda(\raisebox{1.72218pt}{.})=-\log\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$, where we omit the index $D\in\mathbb{N}$ in the above notation. The existence of such $\mathcal{F}_{\mathrm{sp}}$ is ensured by the finiteness of the values of .$\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ on $\overline{E}_{\operatorname{HN},D}$.
@@
By Proposition 2.3.2, the sequence $\left(\frac{1}{n}\widehat{\mu}_{\max}(\overline{E}_{n})\right)_{n\geqslant 0}$ converges to a real number, and we denote it by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\bullet})$. If $E_{n}=H^{0}(X,L^{\otimes n})$ equipped with the superior norms, it is denoted by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$ for simplicity.
@@
where $\lambda_{+}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ is defined at (4) with respect to the canonical filtrated structure introduced in §3.4.3 over $S_{0}$.
@@
We recall some results on the criteria of arithmetic positivities via $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ in [11, §9] below, where we always suppose that $K$ is perfect.
@@ 4.1 Approximatible graded linear series
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $\pi:X\rightarrow\operatorname{Spec}K$ be a projective morphism, and $\overline{L}$ be an adelic line bundle over $S$. With all the notations and assumption same as in (10). If $\overline{L}$ is big, then for all $\lambda\in\left(-\infty,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$, $\overline{E}^{\lambda}_{\operatorname{HN},\bullet}$ contains an ample divisor.
Proof.
@@
[…]
In [4, Proposition 3.11], a similar result of Proposition 4.1.3 is proved by an arithmetic analogue of Kodaira decomposition [25, Corollary 2.4]. Until the author’s knowledge, we do not know whether it is still valid over a general adelic curve. To avoid this obstruction, we use a method quite similar to that of [2], and see also [9, §6.3].
@@ 4.2 Fujita approximation of graded linear series
Let $p\in\mathbb{N}$ such that $E_{\operatorname{HN},p}^{0}\neq 0$, where we follow the notation in (9). In this case, we define $\overline{E}_{\bullet}^{(p)}=\bigoplus\limits_{n\geqslant 0}\overline{E}^{(p)}_{\operatorname{HN},np}$ to be the graded $K$-algebra generated by $E_{\operatorname{HN},p}^{0}$, which is considered as a normed graded linear series over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ defined in Example 2.1.1.
@@
Proof.
For any $n\in\mathbb{N}^{+}$, let $\nu_{\mathrm{sp},n}=T_{\frac{1}{n}}\nu_{\left(E_{\operatorname{HN},n},\mathcal{F}_{\mathrm{sp}}\right)}$, where the notation $T_{\frac{1}{n}}\nu_{\left(E_{\operatorname{HN},n},\mathcal{F}_{\mathrm{sp}}\right)}$ is introduced in §2.3.2, and the $\mathbb{R}$-filtration $\mathcal{F}_{\mathrm{sp}}$ is defined in §3.4.2 with respect to the spectral norms. Then by Proposition 2.3.2, the sequence $(\nu_{\mathrm{sp},n})_{n\geqslant 1}$ converges weakly to a Borel probability measure, which is denoted by $\nu_{\mathrm{sp}}$.
@@
In order to prove the assertion, we will prove that the restriction of $\nu_{\mathrm{sp}}$ on $[0,+\infty)$ can be approximated well by $\nu_{\mathrm{sp}}^{(p)}$ for $p\in\mathbb{N}$.
@@
[…]
be a division of the interval $[0,\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet}))$ such that
@@ 5.1 Fujita approximation of big adelic line bundles
    $\overline{M}$ is effective and $\overline{G}$ is arithmetically nef;
@@
We consider $\overline{E}^{(p)}_{\bullet}$ as a normed graded linear series over $S$, which is equipped with the subspace norms of $\overline{E}_{\bullet}=\bigoplus\limits_{D\geqslant 0}\overline{E}_{D}$. For any $n\geqslant 1$ and $w\in\Omega$, let .$\|\raisebox{1.72218pt}{.}\|_{n,w}$ be the quotient norm on $G_{p,w}=G_{p}\times_{\operatorname{Spec}K}\operatorname{Spec}K_{w}$ induced by the surjective homomorphism
@@
where $E_{\operatorname{HN},np}^{(p)}$ is equipped with the superior norm .$\|\raisebox{1.72218pt}{.}\|_{\sup,w}$ with respect to $w\in\Omega$.
@@
By Hahn-Banach theorem, every linear form on $\ker(E_{\operatorname{HN},np}^{(p)}\rightarrow\ell_{n,x})$ can be extended to a linear form on $E_{\operatorname{HN},np}^{(p)}$. In other words, there exists a section $w\in E_{\operatorname{HN},np}^{(p)}$, whose image in $G_{p,w}^{\otimes n}(x)$ has norm $\|w\|_{\sup,w}$. As a section of $L_{w}^{\otimes np}$ on $X_{v}(\overline{K}_{w})$, we have $\|w(x)\|_{w}\leqslant\|w\|_{\sup,w}$, where .$\|\raisebox{1.72218pt}{.}\|_{w}$ is the metric on $\overline{L}$ indexed by $v\in\Omega$. Then for any section $u\in G_{p,w}$ over a neighborhood of $x$, we may assume that $u(x)^{\otimes n}$ equals to the image of $w(x)$ in $G_{p,w}^{\otimes n}(x)$ by dilation, then we have
@@
Let $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ be a family of adelic metrics on $G_{p}$, and we define $(M_{p},\beta_{n})=\phi_{p}^{*}(\overline{L})\otimes(G_{p},\alpha_{n})^{\vee}$. In this sense, $s$ is an effective section in $(M_{p},\beta_{n})$, then it is an effective adelic line bundle.
@@
and .$\|\raisebox{1.72218pt}{.}\|_{n,\sup,w}$ is the superior norm induced by $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ on $G_{p}$. Then by [4, Corollary A.2], we have
@@
with $\overline{G}$ arithmetically nef and $\overline{M}$ effective, such that
@@
Let $\overline{B}^{\prime}$ be an arithmetically ample line bundle on $X^{\prime}$. Since $\phi^{*}(\overline{B})$ is arithmetically big, there exists a $q\in\mathbb{N}$, such that $\phi^{*}(\overline{B}^{\otimes q})\otimes\overline{B}^{\prime\vee}$ is effective. We consider
@@
In the above decomposition, $\overline{G}^{\prime}:=\overline{G}^{\otimes q}\otimes\overline{B}^{\prime}$ is arithmetically ample, $\overline{M}^{\otimes q}\otimes\phi^{*}(\overline{B}^{\otimes pq})\otimes\overline{B}^{\prime\vee}$ is effective, and

### v3

@@ 3.4 Graded linear series over trivially valued fields
Let $\overline{E}$ be an adelic vector bundle over an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, and $\left(\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E})\right)_{t\in\mathbb{R}}$ be the Harder-Narasimhan filtration of $\overline{E}$ defined in Definition 2.2.3. We define the norm $\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}$ on $E$ over $(K,|\raisebox{1.72218pt}{.}|_{0})$ as
$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}:&E&\longrightarrow&\mathbb{R}^{+}\\ &s&\mapsto&\exp\left(-\sup\{t\in\mathbb{R}\mid s\in\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E})\}\right).\end{array}$ (7)
@@
Let $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ be a normed graded linear series over an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$. For every $\overline{E}_{D}$, we denote by $\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}}$ the norm (7) on $E_{D}$ indexed by $D\in\mathbb{N}$. Then $(E_{D},\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}})$ an adelic vector bundle over $S_{0}$.
@@
$\|s\cdot t\|^{\operatorname{HN}}_{m+n}\leqslant\exp\left(f(\dim E_{m})+f(\dim E_{n})\right)\|s\|^{\operatorname{HN}}_{m}\|t\|^{\operatorname{HN}}_{n},$
@@
In order to obtain a new family of norms which are sub-multiplicative, we introduce the spectral norm $\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ of $\|\raisebox{1.72218pt}{.}\|^{\operatorname{HN}}_{D}$ as
$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}:&H^{0}(X,L^{\otimes D})&\longrightarrow&\mathbb{R}_{\geqslant 0}\\ &s&\mapsto&\lim\limits_{n\rightarrow\infty}\left(\|s^{n}\|_{nD}^{\operatorname{HN}}\right)^{1/n}.\end{array}$ (8)
@@
We denote by $\mathcal{F}_{\mathrm{sp}}$ the $\mathbb{R}$-filtration on $\overline{E}_{\operatorname{HN},D}$ induced by the filtrated structure $\lambda(\raisebox{1.72218pt}{.})=-\log\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$, where we omit the index $D\in\mathbb{N}$ in the above notation. The existence of such $\mathcal{F}_{\mathrm{sp}}$ is ensured by the finiteness of the values of $\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ on $\overline{E}_{\operatorname{HN},D}$.
@@
By Proposition 2.3.2, the sequence $\left(\frac{1}{n}\widehat{\mu}_{\max}(\overline{E}_{n})\right)_{n\geqslant 0}$ converges to a real number, and we denote it by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\bullet})$. If $E_{n}=H^{0}(X,L^{\otimes n})$ equipped with the supremum norms, it is denoted by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$ for simplicity.
@@
[…]
@@
[…]
@@ 4.1 Approximatible graded linear series
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $\pi:X\rightarrow\operatorname{Spec}K$ be a projective morphism, and $\overline{L}$ be an adelic line bundle over $S$. With all the notations and assumption same as in (10). If $\overline{L}$ is big, then for all $\lambda\in\left(-\infty,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$, $\overline{E}^{\lambda}_{\operatorname{HN},\bullet}$ contains an ample divisor.
Démonstration.
@@
4.1.3
In [4, Proposition 3.11], a similar result of Proposition 4.1.2 is proved by an arithmetic analogue of Kodaira decomposition [25, Corollary 2.4] of X. Yuan. But it is not valid over a general adelic curve, and see Example 4.1.3 below for a counterexample. To avoid this obstruction, we use a method quite similar to that of [2] to prove Proposition 4.1.2, and see also [9, §6.3].
[…]
Let $K$ be a field, and $S_{3}=(K,\{1,2,3\},|\raisebox{1.72218pt}{.}|_{0})$, where the trivially valued absolute $|\raisebox{1.72218pt}{.}|_{0}$ (see Example 2.1.1 for the precise definition) is counted three times in a discrete $\sigma$-algebra. We consider the universal line bundle $\mathcal{O}(1)$ on $\mathbb{P}_{K}^{1}$, and we denote by $x_{0},x_{1}$ the basis of $H^{0}(\mathbb{P}^{1}_{K},\mathcal{O}(1))$ with respect to the homogeneous coordinate.
Let $\overline{\mathcal{O}(1)}=(\mathcal{O}(1),(\|\raisebox{1.72218pt}{.}\|_{i})_{i=1}^{3})$ be the adelic line bundle on $\mathbb{P}^{1}_{K}$, where the induced supremum norms $\|\raisebox{1.72218pt}{.}\|_{i=1}^{3}$ on $H^{0}(\mathbb{P}^{1}_{K},\mathcal{O}(1))$ satisfies
$\begin{cases}\|x_{i}\|_{1}=e^{-\frac{1}{2},},\text{ where }i=0,1;\\ \|x_{i}\|_{2}=e^{-\frac{1}{2},},\text{ where }i=0,1;\\ \|x_{i}\|_{3}=e^{\frac{3}{4},},\text{ where }i=0,1.\end{cases}$
Then $\overline{\mathcal{O}(1)}$ is an arithmetically big line bundle on $\mathbb{P}^{1}_{K}$ over $S_{3}$, since
$\widehat{\deg}\left(H^{0}(\mathbb{P}^{1}_{K},\mathcal{O}(D)),(\|\raisebox{1.72218pt}{.}\|_{i})_{i=1}^{3}\right)=\frac{1}{4}D(D+1)$
for all $D\in\mathbb{N}^{+}$.
If there exist a $D\in\mathbb{N}$, an adelic ample line bundle $\overline{A}$ and an line bundle $\overline{M}$ have a non-zero small section on $\mathbb{P}^{1}_{K}$, such that $\overline{\mathcal{O}(D)}=\overline{A}\otimes\overline{M}$, where $\overline{\mathcal{O}(D)}$ is equipped with the induced metrics from the one introduced above. Since $\operatorname{Pic}(\mathbb{P}^{1}_{K})=\mathbb{Z}$, then there exists a $D^{\prime}\in\mathbb{N}$ satisfying $D^{\prime}<D+$, such that
$\overline{\mathcal{O}(D)}=\overline{\mathcal{O}(D^{\prime})}\otimes\overline{\mathcal{O}(D-D^{\prime})},$
where $\overline{\mathcal{O}(D^{\prime})}$ is ample and $\overline{\mathcal{O}(D-D^{\prime})}$ has a non-zero small section. By the definition of arithmetically ample line bundle, all the global sections of $\overline{\mathcal{O}(D^{\prime})}$ should be of the norms smaller than $1$. But by the choice of $\|\raisebox{1.72218pt}{.}\|_{3}$ in $H^{0}(\mathbb{P}^{1}_{K},\mathcal{O}(1))$ over $|\raisebox{1.72218pt}{.}|_{0}$, all the global sections of $\overline{\mathcal{O}(D-D^{\prime})}$ should be equipped with a norm strictly larger than $1$ with respect to the third $\|\raisebox{1.72218pt}{.}\|_{0}$ in $S_{3}$. So there will be no small section in $\overline{\mathcal{O}(D-D^{\prime})}$, which shows there is no arithmetic Kodaira decomposition in this case.
@@ 4.2 Fujita approximation of graded linear series
Let $p\in\mathbb{N}$ such that $\overline{E}_{\operatorname{HN},p}^{0}\neq 0$, where we follow the notation in (9). In this case, we define $\overline{E}_{\bullet}^{(p)}=\bigoplus\limits_{n\geqslant 0}\overline{E}^{(p)}_{\operatorname{HN},np}$ to be the graded $K$-algebra generated by $\overline{E}_{\operatorname{HN},p}^{0}$, which is considered as a normed graded linear series over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ defined in Example 2.1.1.
@@
Démonstration.
For any $n\in\mathbb{N}^{+}$, let $\nu_{\mathrm{sp},n}=T_{\frac{1}{n}}\nu_{\left(E_{\operatorname{HN},n},\mathcal{F}_{\mathrm{sp}}\right)}$, where the notation $T_{\frac{1}{n}}\nu_{\left(E_{\operatorname{HN},n},\mathcal{F}_{\mathrm{sp}}\right)}$ is introduced in §2.3.2, and the $\mathbb{R}$-filtration $\mathcal{F}_{\mathrm{sp}}$ is defined in §3.4.2 with respect to the spectral norms. Then by Proposition 2.3.2 and Corollary 4.1.2, the sequence $(\nu_{\mathrm{sp},n})_{n\geqslant 1}$ converges weakly to a Borel probability measure, which is denoted by $\nu_{\mathrm{sp}}$.
@@
In order to prove the assertion, we will prove that the restriction of $\nu_{\mathrm{sp}}$ on $0,+\infty)$ can be approximated well by $\nu_{\mathrm{sp}}^{(p)}$ for $p\in\mathbb{N}$.
@@
$D:\;0=t_{0}<t_{1}<\cdots<t_{m}<\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet})$
be a division of the interval $0,\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet}))$ such that
@@ 5.1 Fujita approximation of big adelic line bundles
    $\overline{M}$ has a non-zero small section and $\overline{G}$ is arithmetically nef ;
@@
We consider $\overline{E}^{(p)}_{\bullet}$ as a normed graded linear series over $S$, which is equipped with the subspace norms of $\overline{E}_{\bullet}=\bigoplus\limits_{D\geqslant 0}\overline{E}_{D}$. For any $n\geqslant 1$ and $w\in\Omega$, let $\|\raisebox{1.72218pt}{.}\|_{n,w}$ be the quotient norm on $G_{p,w}=G_{p}\times_{\operatorname{Spec}K}\operatorname{Spec}K_{w}$ induced by the surjective homomorphism
@@
where $E_{\operatorname{HN},np}^{(p)}$ is equipped with the superior norm $\|\raisebox{1.72218pt}{.}\|_{\sup,w}$ with respect to $w\in\Omega$.
@@
By Hahn-Banach theorem, every linear form on $\ker(E_{\operatorname{HN},np}^{(p)}\rightarrow\ell_{n,x})$ can be extended to a linear form on $E_{\operatorname{HN},np}^{(p)}$. In other words, there exists a section $w\in E_{\operatorname{HN},np}^{(p)}$, whose image in $G_{p,w}^{\otimes n}(x)$ has norm $\|w\|_{\sup,w}$. As a section of $L_{w}^{\otimes np}$ on $X_{v}(\overline{K}_{w})$, we have $\|w(x)\|_{w}\leqslant\|w\|_{\sup,w}$, where $\|\raisebox{1.72218pt}{.}\|_{w}$ is the metric on $\overline{L}$ indexed by $v\in\Omega$. Then for any section $u\in G_{p,w}$ over a neighborhood of $x$, we may assume that $u(x)^{\otimes n}$ equals to the image of $w(x)$ in $G_{p,w}^{\otimes n}(x)$ by dilation, then we have
@@
Let $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ be a family of adelic metrics on $G_{p}$, and we define $(M_{p},\beta_{n})=\phi_{p}^{*}(\overline{L})\otimes(G_{p},\alpha_{n})^{\vee}$. In this sense, $s$ is an small section in $(M_{p},\beta_{n})$, then it is an adelic line bundle have a non-zero small section.
@@
and $\|\raisebox{1.72218pt}{.}\|_{n,\sup,w}$ is the superior norm induced by $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ on $G_{p}$. Then by [4, Corollary A.2], we have
@@
with $\overline{G}$ arithmetically nef and $\overline{M}$ having a non-zero small section, such that
@@
Let $\overline{B}^{\prime}$ be an arithmetically ample line bundle on $X^{\prime}$. Since $\phi^{*}(\overline{B})$ is arithmetically big, there exists a $q\in\mathbb{N}$, such that $\phi^{*}(\overline{B}^{\otimes q})\otimes\overline{B}^{\prime\vee}$ has a non-zero small section. We consider
@@
In the above decomposition, $\overline{G}^{\prime}:=\overline{G}^{\otimes q}\otimes\overline{B}^{\prime}$ is arithmetically ample, $\overline{M}^{\otimes q}\otimes\phi^{*}(\overline{B}^{\otimes pq})\otimes\overline{B}^{\prime\vee}$ has a non-zero small section, and